Determine the given and compute the appropriate test statistic of the problem below.



Construct the rejection region of the problem below





In a study of television viewing, the mean number of television program they watched during daytime was 7. A survey was conducted on the random sample of 25 households and found that the mean number of television program they watched during daytime was 5 with a standard deviation of 1.5. Test the hypothesis at 10% level of significance.

1
Expert's answer
2022-06-06T08:17:07-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=7"

"H_1:\mu\not=7"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.10," "df=n-1=24" and the critical value for a two-tailed test is "t_c = 1.710882."

The rejection region for this two-tailed test is "R = \{t:|t|>1.710882\}."

The t-statistic is computed as follows:



"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{5-7}{1.5/\sqrt{25}}=-6.6667"


Since it is observed that "|t|=6.6667> 1.710882=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed, "df=24" degrees of freedom, "t=-6.6667" is "p= 0.000001," and since "p= 0.000001<0.10=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 7, at the "\alpha = 0.10" significance level.